OpenStudy (anonymous):

please help! Find the mass flow rate of a water wheel if the power output is 25,000 Watts. Then find the efficiency.

OpenStudy (anonymous):

energy of rotation from wheel conserved to power.. \[W=1/2 \ I \ \omega ^{2}, \ P=W/t\] \[P=1/2 \ k\ M\ R^{2} / t \]

OpenStudy (anonymous):

\[sorry\ P=1/2\ k\ M\ R^{2}\ \omega^{2} / t\]

OpenStudy (anonymous):

\[if \ the\ wheel\ is\ solid\ cylinder, \ k=1/2 \]\[ then\ v= \omega \ R\] subtitute that and you will get the answer..

OpenStudy (anonymous):

i think your question isn't complete..

OpenStudy (anonymous):

hmm thanks, im still a bit unsure of this question. I think something is missing as well but this is what they gave us...